# Positive microlocal holonomies are globally regular

### v1

@@ 2.3. A $\mathfrak{t}$-action on the derived moduli stacks of sheaves
The freedom of the choosing the location of the basepoints $\mathfrak{t}\subset\Lambda$ yields interesting automorphisms of $\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0})$. Though these automorphisms are not strictly needed to prove Theorem 1.1, they provide a conceptually useful perspective on the torus actions studied in [CGGS20, CGG+22] and, in some cases, allow for neater descriptions of the derived stacks $\mathfrak{M}{(\Lambda,\mathfrak{t})}$ and $\mathfrak{M}(\Lambda)$, see Subsection 2.5. These automorphisms are constructed as follows.
Given $\mathfrak{t}$, let us order the components of $\Lambda$ as $\Lambda=\Lambda_{1}\cup\ldots\cup\Lambda_{p}$, $p=|\pi_{0}(\Lambda)|$, and suppose that the basepoints in $\mathfrak{t}$ contained in the $k$-th component are labeled by $t_{k,1},\ldots,t_{k,i_{k}}$. This labeling $t_{k,1},\ldots,t_{k,i_{k}}$ is cyclically ordered, in that we choose it such that $t_{k,j}$ and $t_{k,j+1}$ (and $t_{k,i_{k}}$ and $t_{k,1}$) are consecutive as we traverse the $k$-th component of $\Lambda$. That is, with the given orientation on $\Lambda_{k}$, the oriented path from $t_{k,j}$ to $t_{k,j+1}$ has no $\mathfrak{t}$-basepoints, for all $j\in\mathbb{Z}/i_{k}\mathbb{Z}$.
Consider a family of base points $\{\mathfrak{t}_{s}\}_{s\in[0,1]}$, $\mathfrak{t}_{s}\subset\Lambda$, smoothly varying in $s$. The smooth family extends to an ambient contact isotopy of $T^{*}\mathbb{R}^{2}$ which (set theoretically) fixes the Legendrian $\Lambda\subset T^{*}_{\infty}\mathbb{R}^{2}$. By construction, there is a quasi-equivalence $P(\mathfrak{t}_{s}):\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}_{0}}(\mathbb{R}^{2})_{0}\longrightarrow\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}_{s}}(\mathbb{R}^{2})_{0}$ for any $s\in[0,1]$. There is a particular family of automorphisms that will induce self-equivalences relevant to us, constructed as follows.
Definition 2.3.
A family $\{\mathfrak{t}^{\circlearrowright,k}_{s}\}_{s\in[0,1]}$ of basepoints is said to go around the $k$-th clock if it has the following properties:
@@
    $\mathfrak{t}^{\circlearrowright,k}_{s}\subset\Lambda$ is a family of basepoints, each with at least one basepoint per component, $s\in[0,1]$, and such that $\mathfrak{t}^{\circlearrowright,k}_{0}=\mathfrak{t}^{\circlearrowright,k}_{1}$.
@@
    Each point $t_{k,j}\in\mathfrak{t}_{0}^{\circlearrowright,k}$ in the $k$-th component $\Lambda_{k}\subset\Lambda$ moves from its initial position $t_{k,j}\in\mathfrak{t}_{0}^{\circlearrowright,k}$ to the position of $t_{k,{j+1}}\in\mathfrak{t}_{1}^{\circlearrowright,k}=\mathfrak{t}^{\circlearrowright,k}_{0}$ as $s\in[0,1]$ ranges from $s=0$ to $s=1$ and does not intersect $\mathfrak{t}^{\circlearrowright,k}_{0}$ for $s\in(0,1)$.
  3. (3)
    Each point in $\mathfrak{t}_{s}^{\circlearrowright,j}$ in the $j$-th component of $\Lambda$ stays fixed for all $s\in[0,1]$ if $j\neq k$. $\Box$
Given such a family in Definition 2.3, Property $(1)$ guarantees that the corresponding functor $P(\mathfrak{t}_{1}^{\circlearrowright,k}):\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}_{0}}(\mathbb{R}^{2})\longrightarrow\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}_{1}}(\mathbb{R}^{2})$ is a self-equivalence of $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}_{0}}(\mathbb{R}^{2})$. By Property $(2)$, this self-equivalence will be non-trivial in certain cases. For such a family, we denote the above self-equivalences $P(\mathfrak{t}_{1}^{\circlearrowright,k})$ of $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})=\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}_{0}}(\mathbb{R}^{2})$ by $P^{\circlearrowright,k}$. The induced automorphisms in $\mathfrak{M}({\Lambda,\mathfrak{t}})$ are denoted by $\mathbb{P}^{\circlearrowright,k}$. This construction defines an algebraic action of the free group $F_{p}:=\langle x_{1},\ldots,x_{p}\rangle$ on the $D^{-}$-stack $\mathfrak{M}({\Lambda,\mathfrak{t}})$ by letting the generator $x_{k}\in F_{p}$ act on $\mathfrak{M}({\Lambda,\mathfrak{t}})$ via $\mathbb{P}^{\circlearrowright,k}$. This extends to an action of $F_{p}$ by composing these automorphisms $\mathbb{P}^{\circlearrowright,k}$.
Definition 2.4.
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{st})$ be a Legendrian link and $\mathfrak{t}\subset\Lambda$ a set of basepoints, with at least one basepoint per component of $\Lambda$. By definition, the basepoint action of the free group $F_{p}$ on $\mathfrak{M}({\Lambda,\mathfrak{t}})$ is the action generated by the automorphisms $\mathbb{P}^{\circlearrowright,k}$ described above, $k\in[1,p]$. $\Box$
Example 2.5.
$(i)$ Let $(\Lambda,\mathfrak{t})$ have a unique basepoint in its $k$-th component. Then $\mathbb{P}^{\circlearrowright,k}\in\operatorname{Aut}(\mathfrak{M}({\Lambda,\mathfrak{t}}))$ acts on a point with underlying sheaf $\mathscr{F}\in\mathfrak{M}({\Lambda,\mathfrak{t}})$ by applying the microlocal monodromy of the local system $m_{\Lambda}(\mathscr{F})$ associated to $\mathscr{F}$ (restricted to the $k$-th component of $\Lambda$) to the microstalk at the unique basepoint in the $k$th component. That is, the action of $F_{p}$ factors through the natural monodromy action of the local system $m_{\Lambda}(\mathscr{F})\in\mathrm{Loc}(\Lambda)$ on its stalks. See Lemma 2.6 below where this is justified.
$(ii)$ In the case that the $k$-th component $\Lambda_{k}$ has $i_{k}$ basepoints, the same geometric picture holds for $\mathbb{P}^{\circlearrowright,k}$after considering $i_{k}$-th roots. Namely, the action of $x_{k}$ is via parallel transport according to $m_{\Lambda}(\mathscr{F})$ from one basepoint of $\Lambda_{k}$ to the next. Indeed, since all the stalks are identified – as we pull-backed along the diagonal morphism – this parallel transport yields automorphisms, not just isomorphisms between those stalks. In particular, the $i_{k}$-th power of $\mathbb{P}^{\circlearrowright,k}$applies the monodromy of $m_{\Lambda}(\mathscr{F})$ along $\Lambda_{k}$ to the stalks at the basepoints. $\Box$
The most important property of the action in Definition 2.4, and the reason we present it explicitly, is that it presents $\mathfrak{M}({\Lambda})$ as a quotient of $\mathfrak{M}({\Lambda,\mathfrak{t}})$, as follows from Lemma 2.6 below. Precisely, the around-the-clock action in Definition 2.3 induced by moving around basepoints typically acts non-trivially in $\mathfrak{M}({\Lambda,\mathfrak{t}})$, cf. e.g. Example 2.5, and its quotient yields $\mathfrak{M}({\Lambda})$. Thus we construct $\mathfrak{M}({\Lambda,\mathfrak{t}})$ from $\mathfrak{M}({\Lambda})$ via a homotopy pull-back and recover $\mathfrak{M}({\Lambda})$ from $\mathfrak{M}({\Lambda,\mathfrak{t}})$ via the quotient given by this explicit action.
In order to describe the automorphisms $\mathbb{P}^{\circlearrowright,k}$ of $\mathfrak{M}({\Lambda,\mathfrak{t}})$ in terms of linear algebra, consider the connected component $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$ consisting of sheaves with microlocal rank $n$ and suppose that $\mathfrak{t}$ has exactly one basepoint per component of $\Lambda$, so $|\mathfrak{t}|=p$. Let us describe a point in $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})$ by a pair (,$(\mathscr{F},\{\phi_{i}\})$ consisting of a sheaf $\mathscr{F}$ of microlocal rank $n$ and a set of (common) trivializations $\{\phi_{i}\}_{i\in[1,p]}$ for the microstalks of $\mathscr{F}$ at the $p$ basepoints in $\mathfrak{t}$. The key lemma reads as follows:
Lemma 2.6.
Let $(\Lambda,\mathfrak{t})$ have exactly one basepoint per component. Then the automorphism $\mathbb{P}^{\circlearrowright,k}\in\operatorname{Aut}(\mathfrak{M}({\Lambda,\mathfrak{t}}))$ is given by
$\mathbb{P}^{\circlearrowright,k}((\mathscr{F},\{\phi_{i}\}))=(\mathscr{F},\{\widehat{\phi}_{i}\}),$
with $\widehat{\phi}_{k}:={\mbox{mon}}_{\Lambda_{k}}(\mathscr{F})\circ\phi_{k}$, where ${\mbox{mon}}_{\Lambda_{k}}(\mathscr{F})$ denotes the microlocal monodromy of $m_{\Lambda}(\mathscr{F})$ along the $k$-th component $\Lambda_{k}\subset\Lambda$ and, if otherwise $i\neq k$, $\widehat{\phi}_{i}:=\phi_{i}$.
Proof.
Since $\mathbb{P}^{\circlearrowright,k}\in\operatorname{Aut}(\mathfrak{M}({\Lambda,\mathfrak{t}}))$ preserves $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})$ for any $n\in\mathbb{N}$, we focus on the action on such a connected component and denote $G:=\operatorname{GL}_{n}$. For any local system, the stalk determines a point in $BG$. For any local system on a circle, it determines a map $S^{1}\longrightarrow BG$, and the monodromy, as a function on the loop space, is given by the induced map $\Omega S^{1}\longrightarrow\Omega BG$. When moving a basepoint around a component $\Lambda_{k}\subset\Lambda$, the automorphism on the stalk of the local system is thus given by applying the loop functor $\Omega$ to that $BG$ stalk. In our case, the movement occurs via the around the $k$-th clock automorphism $\mathbb{P}^{\circlearrowright,k}$ and the stalks are those of the local system on $\Lambda$ given by the microlocal functor applied to $\mathscr{F}$. (So they are the microstalks of $\mathscr{F}$ on $\Lambda_{k}$.) Since $\Omega BG\simeq G/G^{ad}$ and $G/G^{ad}$ is the moduli stack of local systems on $\Lambda_{k}$, the claim follows. ∎

### v2

@@ 2.3. A $\mathfrak{t}$-action on the derived moduli stacks of sheaves
The freedom of the choice for the location of the basepoints $\mathfrak{t}\subset\Lambda$ yields interesting automorphisms of $\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0})$. Though these automorphisms are not strictly needed to prove Theorem 1.1, they provide a conceptually useful perspective on the torus actions studied in [CGG+24, CGG+25] and, in some cases, allow for neater descriptions of the derived stacks $\mathfrak{M}{(\Lambda,\mathfrak{t})}$ and $\mathfrak{M}(\Lambda)$, see Subsection 2.5. These automorphisms are constructed as follows.
Given a set of base points $\mathfrak{t}\subset\Lambda$, let us describe an object in $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ by a pair $(\mathscr{F},(\phi_{i}))$ consisting of a sheaf $\mathscr{F}$ and a set of common trivializations $(\phi_{i})_{i\in[1,|\mathfrak{t}|]}$ for the microstalks of $\mathscr{F}$ at the basepoints in $\mathfrak{t}$, where $\phi_{i}:V\longrightarrow m_{\Lambda,t_{i}}(\mathscr{F})$ are quasi-isomorphisms from a given fixed chain complex $V$. Consider the $\infty$-group of automorphisms
$G(\mathfrak{t}):=\operatorname{Aut}\left(\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)\right),$
understood as a group object in $\infty$-groupoids, see e.g. [LUR09, Section 6.1.2] or [GJ99, Chapter V]. Such automorphisms yield automorphisms of $\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}$ that reparametrize the trivialization at the microstalks, as made precise in the following:
Lemma 2.3.
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{st})$ be a Legendrian link and $\mathfrak{t}\subset\Lambda$ a set of basepoints. Then there is a natural morphism of $\infty$-groups
$P:G(\mathfrak{t})\longrightarrow\operatorname{Aut}\left(\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}\right)$
such that $(c_{i})\in G(\mathfrak{t})$ acts on an object $(\mathscr{F},(\phi_{i}))\in\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}$ by
$P_{(c_{i})}(\mathscr{F},(\phi_{i}))=(\mathscr{F},(c_{i}\circ\phi_{i})).$
Proof.
The homotopy colimit diagram (2.2) of dg-categories reads
$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)$$\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0}$$\operatorname{Perf}(k)$$\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}.$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
By construction, $G(\mathfrak{t})$ acts on the upper-left category. The $\infty$-group morphism $P$ is built by extending this action to $\operatorname{Perf}(k)$ and $\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0}$ so that $\Delta^{\ell}$ and $m_{\Lambda,\mathfrak{t}}^{\ell}$ are respectively equivariant, thus inducing an action on the homotopy pushout $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$. For that, we first consider the trivial action of $G(\mathfrak{t})$ on the bottom left dg-category $\operatorname{Perf}(k)$. Now, for any linear map $\phi:V\longrightarrow W$ and $(c_{i})\in G(\mathfrak{t})$, the corresponding map $\phi^{|\mathfrak{t}|}:V^{|\mathfrak{t}|}\longrightarrow W^{|\mathfrak{t}|}$ induced by the diagonal functor is homotopic to its conjugation $(c_{i}^{-1})\circ\phi^{|\mathfrak{t}|}\circ(c_{i}):V^{|\mathfrak{t}|}\longrightarrow W^{|\mathfrak{t}|}$. Therefore, the natural action of $G(\mathfrak{t})$ on the domain of $\Delta^{\ell}$ and the trivial action of $G(\mathfrak{t})$ on its codomain are compatible. Second, similar considerations apply to $m_{\Lambda,\mathfrak{t}}^{\ell}$. Indeed, by construction, this functor factors as
$m_{\Lambda,\mathfrak{t}}^{\ell}:\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)\longrightarrow\mathrm{Loc}^{c}(\Lambda)\longrightarrow\operatorname{Sh}_{\Lambda}^{c}(\mathbb{R}^{2})_{0},$
and we consider the trivial action of $G(\mathfrak{t})$ on $\mathrm{Loc}^{c}(\Lambda)$ and $\operatorname{Sh}_{\Lambda}^{c}(\mathbb{R}^{2})_{0}$. As before, given any morphism of local system $\phi:(V,\mu)\longrightarrow(W,\eta)$ and $(c_{i})\in G(\mathfrak{t})$, the morphism $\phi^{|\mathfrak{t}|}:V^{|\mathfrak{t}|}\longrightarrow W^{|\mathfrak{t}|}$ induced by the restriction functor to $\mathfrak{t}$ is homotopic to the morphism $(c_{i}^{-1})\circ\phi^{|\mathfrak{t}|}\circ(c_{i}):V^{|\mathfrak{t}|}\longrightarrow W^{|\mathfrak{t}|}$. Hence the natural action of $G(\mathfrak{t})$ on the domain of $m_{\Lambda,\mathfrak{t}}^{\ell}$ is compatible with the trivial action of $G(\mathfrak{t})$ on its codomain.
Since the dg-nerves of the three dg-categories involved [^5] Here we use again that $m_{\Lambda,\mathfrak{t}}^{\ell}$ factors through $\mathrm{Loc}^{c}(\Lambda)$. in the homotopy colimit diagram (2.2), i.e.
$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k),\,\operatorname{Perf}(k)\,\mbox{ and }\,\mathrm{Loc}^{c}(\Lambda),$
are all 1-dimensional simplicial sets, the higher coherence data is automatically compatible under the above functors. Therefore, we obtain an action on their homotopy colimit $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$.
The formula for the action on objects follows by construction. Indeed, since an object $(\mathscr{F},(\phi_{i}))\in\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ is determined by a sheaf $\mathscr{F}\in\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ together with a common trivialization $(\phi_{i})$ of its microstalks, via
$m_{\Lambda,\mathfrak{t}}:\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}\longrightarrow\mathrm{Loc}(\mathfrak{t})\xrightarrow{\sim}\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Mod}(k),$
any element $(c_{i})\in G(\mathfrak{t})$ acts as the identity on the sheaf in $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ and acts on the trivialization maps $(\phi_{i})$ by left multiplication, as this is how $G(\mathfrak{t})$ acts componentwise on (products of) perfect modules. ∎
Remark 2.4.
In the proof of Lemma 2.3, we can alternatively verify the compatibility of the higher coherence data of the action as follows. Consider the universal Cartesian filtration over $\mathrm{B}G(\mathfrak{t})$, whose fiber is the dg-nerve of $\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)$ by straightening, see e.g.  [LUR09, Section 3.2.1]. Consider also the trivial Cartesian fibrations over $\mathrm{B}G(\mathfrak{t})$ whose fibers are $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ and $\operatorname{Perf}(k)$, respectively. Then taking the homotopy colimit over the base gives a Cartesian fibration over $\mathrm{B}G(\mathfrak{t})$ whose fiber is $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ and it is equivalent to the group action in Lemma 2.3 by unstraightening, cf.  [LUR09, Section 3.1.3]. $\Box$
Example 2.5.
To illustrate the action of Lemma 2.3 in a simple instance, consider the following homotopy colimit diagram
(2.4) $\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)$$\operatorname{Perf}(k)$$\operatorname{Perf}(k)$$\mathrm{Loc}^{c}\left(\bigvee_{i=1}^{|\mathfrak{t}|-1}S^{1}\right).$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
Here we have replaced $\operatorname{Sh}_{\Lambda}^{c}(\mathbb{R}^{2})_{0}$ by $\operatorname{Perf}(k)$ and $\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}$ by $\mathrm{Loc}^{c}(\bigvee_{i=1}^{|\mathfrak{t}|-1}S^{1})$ in the homotopy colimit (2.2), which is for instance what happens in the case of $\Lambda$ being the max-tb Legendrian unknot. An object in the colimit category of Eq. 2.4 is given by a pair $(U,(\phi_{i}))$, where $U\in\operatorname{Perf}(k)$ and $\phi_{i}:U\longrightarrow V$ are isomorphisms with a given complex $V$. Such data determines the monodromy around the $i$th circle $S^{1}$ in the wedge, which is given by $\phi_{i+1}\circ\phi_{i}^{-1}:V\longrightarrow V$. Therefore, the action $(V,(\phi_{i}))\longrightarrow(V,(c_{i}\circ\phi_{i}))$ in Lemma 2.3 modifies the monodromy by conjugation, i.e. via
$(\phi_{i+1}\circ\phi_{i}^{-1})\longmapsto c_{i+1}\circ(\phi_{i+1}\circ\phi_{i}^{-1})\circ c_{i}^{-1}.$
We denote the action of $(c_{i})_{i\in[1,|\mathfrak{t}|]}\in G(\mathfrak{t})$ on $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ via Lemma 2.3 by
$P_{(c_{i})}:\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}\longrightarrow\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0},$
and the induced action on derived the moduli stack $\mathfrak{M}(\Lambda,\mathfrak{t})$ by
$\mathbb{P}_{(c_{i})}:\mathfrak{M}(\Lambda,\mathfrak{t})\longrightarrow\mathfrak{M}(\Lambda,\mathfrak{t}).$
Note that the $G(\mathfrak{t})$-action on $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ preserves microstalks. Hence, when we restrict to the connected component $\mathfrak{M}_{n}(\Lambda,\mathfrak{t})\subset\mathfrak{M}(\Lambda,\mathfrak{t})$ consisting of microlocal rank $n$ sheaves, the $G(\mathfrak{t})$-action above yields an algebraic action of the algebraic group $\prod_{i=1}^{|\mathfrak{t}|}\operatorname{GL}_{n}$ on the derived stack $\mathfrak{M}_{n}(\Lambda,\mathfrak{t})$.
Example 2.6.
Let $(\Lambda,\mathfrak{t})$ have a unique basepoint in its $k$-th component. Given an automorphism $c_{k}\in\operatorname{Aut}(\operatorname{Perf}(k))$, the induced automorphism $\mathbb{P}_{c_{k}}\in\operatorname{Aut}(\mathfrak{M}({\Lambda,\mathfrak{t}}))$ acts on a point $(\mathscr{F},(\phi_{i}))$, with underlying sheaf $\mathscr{F}\in\mathfrak{M}({\Lambda})$, by left multiplication $c_{k}\circ\phi_{k}$ on the trivialization $\phi_{k}$ of the microstalk at that unique basepoint in the $k$-th component. This action factors through the natural action on the local system $m_{\Lambda}(\mathscr{F})\in\mathrm{Loc}(\Lambda)$ and, as in Example 2.5, it modifies the microlocal monodromy $mon(\mathscr{F})_{k}$ at the basepoint by conjugation, i.e.  $mon(\mathscr{F})_{k}\longmapsto c_{k}\circ mon(\mathscr{F})_{k}\circ c_{k}^{-1}$. $\Box$
The most important property of the action in Lemma 2.3, and the reason we present it explicitly, is that it presents the stack $\mathfrak{M}_{n}({\Lambda})$ as a quotient of $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})$. Indeed, by the homotopy colimit diagram (2.2), we have a homotopy pull-back diagram
$\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})$$\mathfrak{M}_{n}({\Lambda})$$\operatorname{BGL}_{n}$$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{BGL}_{n}.$}
By construction, the bottom functor between classifying spaces is induced by the diagonal inclusion $\operatorname{GL}_{n}\longrightarrow\prod_{i=1}^{|\mathfrak{t}|}\operatorname{GL}_{n}$, and thus can be presented as a homotopy quotient by $\prod_{i=1}^{|\mathfrak{t}|-1}\operatorname{GL}_{n}$. Therefore, the stack $\mathfrak{M}_{n}({\Lambda})$ can be constructed from $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})$ by quotienting with $\prod_{i=1}^{|\mathfrak{t}|-1}\operatorname{GL}_{n}$ via the action from Lemma 2.3.
Remark 2.7 (A comparison with the basepoint action in the Floer-theoretic perspective).
The action in Lemma 2.3 defines a natural transformation of the identity functor $\mbox{id}:\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}\longrightarrow\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}$ which acts on the trivializations by left multiplication by the monodromy around that component. Such action can be understood by taking based loops on the morphism $P$ of Lemma 2.3. In particular, for microlocal rank 1, we have morphisms
$\Omega\operatorname{GL}_{1}\longrightarrow\Omega_{\mathrm{id}}\operatorname{Aut}(\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0})\xrightarrow{\sim}\operatorname{Aut}(\mathrm{id}_{\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}}).$
Since $\Omega\operatorname{GL}_{1}\cong\mathbb{Z}$, we have a natural automorphism $t_{\mathscr{F}}$ for any $\mathscr{F}\in\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}$. This automorphism can be described geometrically as follows.
Consider a family of base points $\{\mathfrak{t}_{s}\}_{s\in[0,1]}$, $\mathfrak{t}_{s}\subset\Lambda$, varying smoothly on the parameter $s\in[0,1]$. Let us say that such family goes around the $k$-th clock if $\mathfrak{t}_{0}=\mathfrak{t}_{1}$ and
@@
    each point $t_{k}\in\mathfrak{t}_{0}$ in the $k$-th component $\Lambda_{k}\subset\Lambda$ moves around once from $s=0$ to $s=1$ and does not intersect $\mathfrak{t}_{0}$ for $s\in(0,1)$,
@@
    each point in $\mathfrak{t}_{s}$ in the $j$-th component stays fixed for all $s\in[0,1]$ if $j\neq k$.
Then the above automorphism $t_{\mathscr{F}}$ is componentwise given by moving the basepoint around the $k$th clock, which changes the trivialization of the microstalk by left multiplication of the monodromy around that $k$th component. This coincides with the Floer-theoretic description in [CN22, Section 5], which presents such actions from the perspective of the Legendrian contact dg-algebra of $(\Lambda,\mathfrak{t})$. $\Box$
2.3.1. Multiple basepoints per component